OpenStudy (anonymous):

The density of an object is defined as mass per unit volume. Mathematically we write that: D = M/V. Rearrange this expression for: (a) M = ? (b) V = ?

OpenStudy (aaronq):

if you have x=y/z you can solve for y by doing the inverse process of z y=x*z do the same thing

OpenStudy (souvik):

M=D*V V=M/D